OpenStudy (anonymous):

Just invests an amount at 4% interest compounded annually after three years he has $1687.30 what was the original amount Jeff invested? $1200 $1450 $1500 $750

OpenStudy (anonymous):

Okay just work backwards for this question. At the end of 3 years - $1687.30 At the end of 2 years - (100 / 104) x $1687.30 = $1622.40 (to 2dp) At the end of 1 year - (100 / 104) x $1622.40 = $1560 At the start - (100 / 104) x $1560 = $1500

OpenStudy (anonymous):

Thank you so much❤️